A solution contains 19 conjugate acid base pairs. One of these is piperazine and its conjugate acid with a [A-]/[HA] ration of 0.64. What is the pH of the solution? What can you say about the other 18 conjugate pairs?
Answer #1

Mixture of piperazine and its conjugate acid produces a basic buffer.

pOH of the buffer is:

pOH = pKb + log [ conjugate acid]/ [base]

pOH = 4.19 + log (1/0.64) = 4.38

pH = 14 - 4.38 = 9.62

Other 18 conjugate pairs do not form buffers.